# ISWC ’24, Special Session on LLMs, Slot 2: Bridging the Gap: Building Knowledge-Enhanced Gen AI, 11-15 Nov 2024

### v2

@@ 2 System Design
Figure 1: LLM-based SPARQL Query Generator System Architecture
@@ 2.1 Generating embeddings and indexing
Additionally, we generate and index shapes for the classes present in each endpoint. For this, we first retrieve the VoID descriptions from each endpoint, which detail the relationships between subject classes and object classes or datatypes via specific predicates. For example, the Protein class is linked to the Gene class through the encodedBy predicate. The VoID generator is available open source666github.com/JervenBolleman/void-generator and can be used to generate statistics over any SPARQL endpoint. This information allows us to generate simple, human-readable ShEx shapes for each class. Additionally, each object property references a list of classes rather than another shape, making each shape self-contained and interpretable on its own.
@@ 2.2 Prompt building
When a user asks a question, embeddings are generated for this question and the most similar questions are retrieved from the vector database, as well as the most similar classes labels to this question. The questions and their associated query, and the classes label and their associated schema are added to the prompt alongside the user question. An example of the generated prompt can be found online 777github.com/sib-swiss/sparql-llm/blob/main/notebooks/EXAMPLE_PROMPT.md.
@@ 2.3 Validating generated queries
To mitigate potential errors (e.g., hallucinations), we developed a method to validate federated SPARQL queries based on the VoID description of the endpoints. The validator parses the SPARQL query generated, extracts triple patterns, identifies the endpoints where they will be executed, and whether the triples in the query comply with the expected schema.
@@
2.4 Logs and feedback
All user questions are stored in a log file, from which they can be analysed to better understand users needs. Additionally, the chat interface includes a simple user feedback mechanism through like/dislike buttons. When one of these buttons is clicked, the entire conversation is stored in a log file in JSONL format on the server. We provide a notebook to display all conversations from the logs in a human-readable way to help maintainers analyze feedbacks.
@@ 3 Implementation
The different components of the system are modular, published as a python package888pypi.org/project/sparql-llm, and can be reused in other systems. The main components included are: 1) A module to automatically generate a human-readable ShEx schema for each class extracted from the VoID description of an endpoint, and generate documents to load in a vector database, compatible with the LangChain framework; 2) A module to automatically retrieve query examples from an endpoint and generate documents to load in a vector database, compatible with the LangChain framework; 3) A module to automatically validate a federated SPARQL query based on each endpoint VoID description, providing human-readable error messages and suggested corrections.
@@ 4 Evaluation
So far, our primary focus has been on developing a production system to assist users of the SIB SPARQL endpoints in formulating effective queries. In the future, we aim to conduct a more comprehensive evaluation using standardised benchmarks[11][12] and benchmarking frameworks[2]. However evaluating query results in bioinformatics is often challenging, even for human experts, due to the inherent complexity of both the domain-specific questions and the KGs models. These complexities can introduce a degree of interpretative flexibility when assessing the accuracy or relevance of the results.
@@ Acknowledgements.
We acknowledge contributions and fruitful discussions with the SIB Semantic Web Focus Group, as well as funding provided through the Biodata Resources Team at SIB. This work was also supported by the CHIST-ERA grant CHIST-ERA-22-ORD-09 by SNF grant 20CH21_217482. UniProt is supported by the National Eye Institute (NEI), National Human Genome Research Institute (NHGRI), National Heart, Lung, and Blood Institute (NHLBI), National Institute on Aging (NIA), National Institute of Allergy and Infectious Diseases (NIAID), National Institute of Diabetes and Digestive and Kidney Diseases (NIDDK), National Institute of General Medical Sciences (NIGMS), National Institute of Mental Health (NIMH), and National Cancer Institute (NCI) of the National Institutes of Health (NIH) under grant U24HG007822. UniProt activities at the SIB are additionally supported by the Swiss Federal Government through the State Secretariat for Education, Research and Innovation SERI.

### v3

@@ 2 System Design
Figure 1: LLM-based SPARQL Query Generator System Architecture.
@@ 2.1 Generating embeddings and indexing
Additionally, we generate and index shapes for the classes present in each endpoint. For this, we first retrieve the VoID descriptions from each endpoint, which detail the relationships between subject classes and object classes or datatypes via specific predicates. For example, the Protein class is linked to the Gene class through the encodedBy predicate. The VoID generator is available open source666github.com/JervenBolleman/void-generator and can be used to generate statistics over any SPARQL endpoint. This information allows us to generate simple, human-readable Shape Expressions (ShEx) for each class. Each object property references a list of classes rather than another shape, making each shape self-contained and interpretable on its own.
We chose to use the ShEx schema language instead of the Web Ontology Language (OWL) for several reasons. First, endpoints may not implement all properties and classes described in the OWL ontologies they utilize, as ontologies operate under the open-world assumption. As a consequence, a more precise schema, based on the actual contents of the endpoint, must be generated for improved accuracy. Additionally, ShEx is specifically designed to express schema constraints, and its syntax is closer to the SPARQL syntax, making it more suitable for defining the structure of data in our context. In contrast, OWL requires separate classes to define properties with specified domains and ranges, resulting in a more complex and verbose representation. As a side effect, the increased verbosity would also incur higher costs due to the increase in number of tokens.
@@ 2.2 Prompt building
When a user asks a question, embeddings are generated for the question, and the 20 most similar questions and 15 closest class labels are retrieved from the vector database. A limit is used rather than a similarity threshold to ensure the model has access to enough examples to determine whether the question can be answered based on the available endpoint data. Moreover, even if the retrieved examples have a low similarity score, they still provide hints on how to write the query, and can be useful when combined with the classes schemas. The retrieved questions and their associated queries, as well as the classes label and their associated schema, are added to the prompt alongside the user question. A concrete example of a generated prompt can be found online 777github.com/sib-swiss/sparql-llm/blob/main/notebooks/EXAMPLE_PROMPT.md.
@@ 2.3 Validating generated queries
To mitigate potential errors (e.g., hallucinations), we developed a method to validate federated SPARQL queries based on the VoID description of the endpoints. First, the validator parses the SPARQL query generated, which can be a federated query. Next, the validator extracts triple patterns, identifies the endpoint(s) where these will be executed, and whether the triple patterns comply with the expected schema. For each subject that has an explicit class assigned, the system will check if the stated predicates are available for this class according to the precomputed ShEx schema. The system will also try to infer the classes of other connected entities that do not have a class explicitly defined.
@@
2.4 Context transparency and feedback
The "See relevant references" button in the chat interface allows users to browse the retrieved queries and classes provided as context to the LLM, along with their similarity scores. This feature enables users to investigate references and understand the reasons behind any incorrect generated query.
All user questions are stored in a log file. Additionally, the chat interface includes a simple user feedback mechanism through like/dislike buttons. When one of these buttons is clicked, the entire conversation is stored in a file on the server. A notebook is provided to display all logged conversations in a human-readable way to help maintainers analyze feedbacks.
@@ 3 Implementation
The different components of the system are modular, published as a python package888pypi.org/project/sparql-llm, and can be reused in other systems. The main components included are: 1) A module to automatically generate a human-readable ShEx schema for each class extracted from the VoID description of an endpoint, and generate documents to load in a vector database, compatible with the LangChain framework; 2) A module to automatically retrieve query examples from an endpoint and generate documents to load in a vector database; 3) A module to automatically validate a federated SPARQL query based on each endpoint VoID description, providing human-readable error messages and suggested corrections.
@@ 4 Evaluation and discussion
So far, our primary focus has been on developing a production system to assist users of the SIB SPARQL endpoints in formulating effective queries. In the future, we aim to conduct a more comprehensive evaluation using standardised benchmarks[11][12] and benchmarking frameworks[2] to assess the system’s precision, usability, and overall performance across a wider range of queries.
We observed that when the system retrieves context that does not fully align with the user’s question, it typically provides an answer that, while imprecise, helps guide the user towards understanding how they might answer the question using available resources. For example, when asked, "What is the function of the esophagus?" the system may return a query that lists anatomical entities with descriptions where "esophagus" appears in the name. Though not directly answering the user’s question, this query still brings the user closer to a solution by suggesting relevant resources. If the retrieved context is insufficiently relevant, the system explicitly states that it cannot answer the question.
Evaluating query results in bioinformatics is particularly challenging, even for human experts, due to the inherent complexity of both the domain-specific questions and the knowledge graph models. These challenges often introduce interpretative flexibility when determining the accuracy or relevance of the system’s outputs. As such, we expect users—typically researchers—to apply a critical approach when assessing the answers provided by the system.
The main weakness we identified in the current system is the tendency to hallucinate entity identifiers that cannot be found in examples, as well as its reliance on inefficient string comparisons that are not well optimized in SPARQL engines. To address these issues, we plan to introduce entity indexing from the endpoints, along with few-shot prompting, to enhance entity extraction and resolution.
@@ Acknowledgements.
We acknowledge contributions and fruitful discussions with the SIB Semantic Web Focus Group, as well as funding provided through the Biodata Resources Team at SIB. This work was also supported by the CHIST-ERA grant CHIST-ERA-22-ORD-09 by SNF grant 20CH21_217482 and the Swiss Open Research Data Grants (CHORD) in Open Science I, a program coordinated by swissuniversities, grant id “Swiss DBGI-KM”. UniProt is supported by the National Eye Institute (NEI), National Human Genome Research Institute (NHGRI), National Heart, Lung, and Blood Institute (NHLBI), National Institute on Aging (NIA), National Institute of Allergy and Infectious Diseases (NIAID), National Institute of Diabetes and Digestive and Kidney Diseases (NIDDK), National Institute of General Medical Sciences (NIGMS), National Institute of Mental Health (NIMH), and National Cancer Institute (NCI) of the National Institutes of Health (NIH) under grant U24HG007822. UniProt activities at the SIB are additionally supported by the Swiss Federal Government through the State Secretariat for Education, Research and Innovation SERI.